# Crystallization of discrete $N$-particle systems at high temperature

### v1

@@ 8.1. The Plancherel case
Next, since $n\geq m\geq 2$ and $m<|z|+2$, we have $n^{m}\leq(2m)^{n}<(2|z|+4)^{n}$. Moreover, if we take $r$ in (75) large enough so that $r>4$, then we want to prove Eq.˜75 only for those $z$ with $|z|>4$. In this case, we can use $2|z|+4\leq 3|z|$. It follows that $n^{m}<(2|z|+4)^{n}\leq(3|z|)^{n}$. Combining with (76) and (77) then leads to
@@
for some constants $C,C^{\prime}>0$. From the asymptotic expansion of the inverse gamma function, we obtain $\left|\Gamma(z)^{-1}\right|\leq\exp\left(C^{\prime\prime}|z|^{2}\right)$, for some constant $C^{\prime\prime}>0$, as long as $|z|$ is large enough. As a result, $\big|F^{\text{planch}}(z)\big|=\left|\Gamma(z)^{-1}\right|\cdot\big|{}_{1}F_{1}\left(\gamma;z;-\eta\right)\big|\leq C\cdot\exp\big(C^{\prime}|z|+C^{\prime\prime}|z|^{2}\big)$, and therefore Eq.˜75 follows, e.g. when $A=3$, if $r>0$ is large enough.
@@
for some constants $C,C^{\prime}>0$. We can also bound $\left|(z+m-2)^{-1}\cdot\Gamma(z)^{-1}\right|\leq\exp\left(C^{\prime\prime}|z|^{2}\right)$, for some $C^{\prime\prime}>0$, by Cauchy’s integral formula. Combining these two inequalities again gives the desired Eq.˜75 for $A=3$ and large enough $r>0$.
@@ 8.2. The alpha case
Together with the bound $\left|\Gamma(z)^{-1}\right|\leq\exp\left(C^{\prime\prime}|z|^{2}\right)$, already used in the Plancherel case, we obtain the desired Eq.˜78 for $A=3$ and large enough $r>0$.
@@
Let us take $r$ in Eq.˜78 to satisfy $r>\max\{\eta,4\}$, so that we can use $\eta<|z|$; also, $4<|z|$ implies $3|z|\leq e^{|z|-1}$. As a result,
@@
If $|z+m-1|\geq\frac{1}{10}$, then $\left|\frac{(z-\eta)^{\uparrow n}}{z^{\uparrow n}}\right|\lesssim\exp(|z|^{2})$, for all $n\geq m$. Some additional work is needed in the case $n<m$, but the same holds, therefore $\big|{}_{2}F_{1}\left(\gamma,z-\eta;z;c\right)\big|\leq\sum_{n\geq 0}{\left|\frac{(z-\eta)^{\uparrow n}}{z^{\uparrow n}}\right|\cdot\frac{\gamma^{\uparrow n}c^{n}}{n!}}\lesssim\exp(|z|^{2})$. Together with $\left|\Gamma(z)^{-1}\right|\leq\exp\left(C^{\prime\prime}|z|^{2}\right)$, we obtain the desired Eq.˜78 for $A=3$ and large enough $r>0$.
If $|z+m-1|<\frac{1}{10}$, then we can prove that $|(z+m-1)\cdot{}_{2}F_{1}\left(\gamma,z-\eta;z;c\right)|\lesssim\exp(|z|^{2})$ as above, and $\big|(z+m-1)^{-1}\cdot\Gamma(z)^{-1}\big|\leq\exp\big(C^{\prime\prime}|z|^{2}\big)$ by Cauchy’s integral formula, as we did in the Plancherel case. The desired Eq.˜78 follows and the proof is finished.

### v2

@@ 8.1. The Plancherel case
Next, since $n\geq m\geq 2$ and $m<|z|+2$, we have $n^{m}\leq(2m)^{n}<(2|z|+4)^{n}$. Moreover, if we take $r$ in (75) large enough so that $r>4$, then we want to prove Equation˜75 only for those $z$ with $|z|>4$. In this case, we can use $2|z|+4\leq 3|z|$. It follows that $n^{m}<(2|z|+4)^{n}\leq(3|z|)^{n}$. Combining with (76) and (77) then leads to
@@
for some constants $C,C^{\prime}>0$. From the asymptotic expansion of the inverse gamma function, we obtain $\left|\Gamma(z)^{-1}\right|\leq\exp\left(C^{\prime\prime}|z|^{2}\right)$, for some constant $C^{\prime\prime}>0$, as long as $|z|$ is large enough. As a result, $\big|F^{\text{planch}}(z)\big|=\left|\Gamma(z)^{-1}\right|\cdot\big|{}_{1}F_{1}\left(\gamma;z;-\eta\right)\big|\leq C\cdot\exp\big(C^{\prime}|z|+C^{\prime\prime}|z|^{2}\big)$, and therefore Equation˜75 follows, e.g. when $A=3$, if $r>0$ is large enough.
@@
for some constants $C,C^{\prime}>0$. We can also bound $\left|(z+m-2)^{-1}\cdot\Gamma(z)^{-1}\right|\leq\exp\left(C^{\prime\prime}|z|^{2}\right)$, for some $C^{\prime\prime}>0$, by Cauchy’s integral formula. Combining these two inequalities again gives the desired Equation˜75 for $A=3$ and large enough $r>0$.
@@ 8.2. The alpha case
Together with the bound $\left|\Gamma(z)^{-1}\right|\leq\exp\left(C^{\prime\prime}|z|^{2}\right)$, already used in the Plancherel case, we obtain the desired Equation˜78 for $A=3$ and large enough $r>0$.
@@
Let us take $r$ in Equation˜78 to satisfy $r>\max\{\eta,4\}$, so that we can use $\eta<|z|$; also, $4<|z|$ implies $3|z|\leq e^{|z|-1}$. As a result,
@@
If $|z+m-1|\geq\frac{1}{10}$, then $\left|\frac{(z-\eta)^{\uparrow n}}{z^{\uparrow n}}\right|\lesssim\exp(|z|^{2})$, for all $n\geq m$. Some additional work is needed in the case $n<m$, but the same holds, therefore $\big|{}_{2}F_{1}\left(\gamma,z-\eta;z;c\right)\big|\leq\sum_{n\geq 0}{\left|\frac{(z-\eta)^{\uparrow n}}{z^{\uparrow n}}\right|\cdot\frac{\gamma^{\uparrow n}c^{n}}{n!}}\lesssim\exp(|z|^{2})$. Together with $\left|\Gamma(z)^{-1}\right|\leq\exp\left(C^{\prime\prime}|z|^{2}\right)$, we obtain the desired Equation˜78 for $A=3$ and large enough $r>0$.
If $|z+m-1|<\frac{1}{10}$, then we can prove that $|(z+m-1)\cdot{}_{2}F_{1}\left(\gamma,z-\eta;z;c\right)|\lesssim\exp(|z|^{2})$ as above, and $\big|(z+m-1)^{-1}\cdot\Gamma(z)^{-1}\big|\leq\exp\big(C^{\prime\prime}|z|^{2}\big)$ by Cauchy’s integral formula, as we did in the Plancherel case. The desired Equation˜78 follows and the proof is finished.